# Tenspiler: A Verified Lifting-Based Compiler for Tensor Operations

### v1

@@ Appendix B Benchmark Name Mapping
  1. 1.
    color_burn
  2. 2.
    color_dodge
  3. 3.
    darken_blend
  4. 4.
    dissolve_blend
  5. 5.
    lighten_blend
  6. 6.
    linear_burn
  7. 7.
    linear_dodge
  8. 8.
    multiply_blend
  9. 9.
    normal_blend
  10. 10.
    normal_blend_f
  11. 11.
    overlay_blend
  12. 12.
    screen_blend
@@
  1. 1.
    matmul
  2. 2.
    rmsnorm_part1
  3. 3.
    rmsnorm_part2
  4. 4.
    softmax_part1
  5. 5.
    softmax_part2
  6. 6.
    softmax_part3
  7. 7.
    softmax_part4
  8. 8.
    transformer_part1
  9. 9.
    transformer_part2
  10. 10.
    transformer_part3
  11. 11.
    transformer_part4
@@
  1. 1.
    dot
  2. 2.
    gemv
Darknet
  1. 1.
    mag_array
  2. 2.
    matrix_add_matrix
  3. 3.
    mse_array
  4. 4.
    mult_add_into_cpu
  5. 5.
    ol_l2_cpu1
  6. 6.
    ol_l2_cpu2
  7. 7.
    scale_array
  8. 8.
    scale_matrix
  9. 9.
    sum_array
  10. 10.
    translate_array
@@
  1. 1.
    fir_small
  2. 2.
    lmsfir1
  3. 3.
    lmsfir2
DSP
  1. 1.
    matadd
  2. 2.
    matscal
  3. 3.
    matsub
  4. 4.
    vadd
  5. 5.
    vcopy
  6. 6.
    vmul
  7. 7.
    vneg
  8. 8.
    voffset
  9. 9.
    vrecip
  10. 10.
    vscal
  11. 11.
    vsub
  12. 12.
    wvec
Dspstone
  1. 1.
    mat1x3
  2. 2.
    n_real_updates
Makespeare
  1. 1.
    sum_of_squares
Mathfu
  1. 1.
    diveq
  2. 2.
    diveq_sca
  3. 3.
    len
  4. 4.
    len_sq
  5. 5.
    matmul_sca
  6. 6.
    muleq
  7. 7.
    muleq_sca
  8. 8.
    negate
  9. 9.
    pluseq
  10. 10.
    subeq
  11. 11.
    subeq_sca
@@
  1. 1.
    array_inc
  2. 2.
    array_sum
  3. 3.
    cube_in_place
  4. 4.
    fourth_in_place
  5. 5.
    sum_elts
Appendix C Configurations of Hardware Devices
@@
Table 2: Configurations of hardware used for experiments101010 Notices & Disclaimers Performance varies by use, configuration and other factors. Learn more on the Performance Index site. Performance results are based on testing as of dates shown in configurations and may not reflect all publicly available updates. See backup for configuration details. No product or component can be absolutely secure. Intel technologies may require enabled hardware, software or service activation. © Intel Corporation. Intel, the Intel logo, and other Intel marks are trademarks of Intel Corporation or its subsidiaries. Other names and brands may be claimed as the property of others.

### v3

@@ 10 Benchmark Name Mapping
# Blend
1 color_burn
2 color_dodge
3 darken_blend
4 dissolve_blend
5 lighten_blend
6 linear_burn
7 linear_dodge
8 multiply_blend
9 normal_blend
10 normal_blend_f
11 overlay_blend
12 screen_blend
@@
# Llama
1 matmul
2 rmsnorm_part1
3 rmsnorm_part2
4 softmax_part1
5 softmax_part2
6 softmax_part3
7 softmax_part4
8 transformer_part1
9 transformer_part2
10 transformer_part3
11 transformer_part4
Darknet
# Darknet
1 mag_array
2 matrix_add_matrix
3 mse_array
4 mult_add_into_cpu
5 ol_l2_cpu1
6 ol_l2_cpu2
7 scale_array
8 scale_matrix
9 sum_array
10 translate_array
DSP
# DSP
1 matadd
2 matscal
3 matsub
4 vadd
5 vcopy
6 vmul
7 vneg
8 voffset
9 vrecip
10 vscal
11 vsub
12 wvec
Dspstone
# Dspstone
1 mat1x3
2 n_real_updates
Makespeare
# Makespeare
1 sum_of_squares
Mathfu
# Mathfu
1 diveq
2 diveq_sca
3 len
4 len_sq
5 matmul_sca
6 muleq
7 muleq_sca
8 negate
9 pluseq
10 subeq
11 subeq_sca
@@
# Blas
1 dot
2 gemv
@@
# Utdsp
1 fir_small
2 lmsfir1
3 lmsfir2
@@
# Simpl Array
1 array_inc
2 array_sum
3 cube_in_place
4 fourth_in_place
5 sum_elts
11 Configurations of Hardware Devices
@@
Table 1: Configurations of hardware used for experiments.131313 Notices & Disclaimers Performance varies by use, configuration and other factors. Learn more on the Performance Index site. Performance results are based on testing as of dates shown in configurations and may not reflect all publicly available updates. See backup for configuration details. No product or component can be absolutely secure. Intel technologies may require enabled hardware, software or service activation. © Intel Corporation. Intel, the Intel logo, and other Intel marks are trademarks of Intel Corporation or its subsidiaries. Other names and brands may be claimed as the property of others.
12 matmul PTX assembly generated by Numba [^14] This is a reduced version with the computation logic for clarity.
1//
2// Generated by NVIDIA NVVM Compiler
3//
4// Compiler Build ID: CL-31833905
5// Cuda compilation tools, release 11.8, V11.8.89
6// Based on NVVM 7.0.1
7//
8
9.version 7.8
10.target sm_89
11.address_size 64
12
13...
14
15  $L_{_{B}}B0_{9}:mov.u64shr.s64and.b64add.s64mul.lo.s64setp.lt.u64mov.f64mov.u64mov.u64mov.u64@$\$$L_{_{B}}B0_{1}2;\par mov.u64mov.u64mov.u64mov.u64\par$\$$L_{_{B}}B0_{1}1:shr.s64and.b64add.s64add.s64shl.b64add.s64and.b64add.s64shl.b64add.s64ld.global.f32ld.global.f32mul.f32cvt.f64.f32add.f64setp.gt.s64selp.u64add.s64selp.b64add.s64selp.b64shr.s64and.b64add.s64add.s64shl.b64add.s64and.b64add.s64shl.b64add.s64ld.global.f32ld.global.f32mul.f32cvt.f64.f32add.f64setp.gt.s64selp.u64add.s64selp.b64add.s64selp.b64shr.s64and.b64add.s64add.s64shl.b64add.s64and.b64add.s64shl.b64add.s64ld.global.f32ld.global.f32mul.f32cvt.f64.f32add.f64setp.gt.s64selp.u64add.s64selp.b64add.s64selp.b64shr.s64and.b64add.s64add.s64shl.b64add.s64and.b64add.s64shl.b64add.s64ld.global.f32ld.global.f32mul.f32cvt.f64.f32add.f64setp.gt.s64selp.u64add.s64selp.b64add.s64selp.b64add.s64setp.ne.s64@$\$$L_{_{B}}B0_{1}1;\par$\$$L_{_{B}}B0_{1}2:setp.eq.s64@$\$$L_{_{B}}B0_{1}6;\par setp.eq.s64shr.s64and.b64add.s64add.s64shl.b64add.s64and.b64add.s64shl.b64add.s64ld.global.f32ld.global.f32mul.f32cvt.f64.f32add.f64@$\$$L_{_{B}}B0_{1}6;\par setp.gt.s64selp.b64setp.eq.s64shr.s64and.b64add.s64add.s64shl.b64add.s64and.b64add.s64shl.b64add.s64ld.global.f32ld.global.f32mul.f32cvt.f64.f32add.f64selp.b64add.s64setp.gt.s64selp.u64add.s64selp.b64@$\$$L_{_{B}}B0_{1}6;\par shr.s64and.b64add.s64add.s64shl.b64add.s64and.b64add.s64shl.b64add.s64ld.global.f32ld.global.f32mul.f32cvt.f64.f32add.f64\par$\$$L_{_{B}}B0_{1}6:and.b64add.s64shl.b64add.s64cvt.rn.f32.f64st.global.f32[setp.gt.s64selp.u64add.s64selp.b64add.s64selp.b64@$\$$L_{_{B}}B0_{9};bra.uni$\$
L__BB0_17;
16
17...
18
19}